# A weakly-modeled search for compact binary coalescences in Einstein Telescope

### v1

@@ IV Effect of the CBC foreground on the detectability of other transient signals
An important issue regarding ET data analysis is the effect of the CBC foreground on the detection capabilities of other GW sources that are expected to be present in the data. In the case of stochastic GW backgrounds, it has been shown that the CBC foreground will indeed prevent the observation of a stochastic background of cosmological origin studies, which motivates the development of methods to accurately substract CBC signals from the data. Here, we study how CBC events recovered by PySTAMPAS affect the ability of the pipeline to detect other kind of unmodelled signals. In order to do that, we inject a selection of waveforms routinely used to characterize unmodelled long-duration GW searches on top of the MDC data, and on top of pure Gaussian noise following the same sensitivity curve as the MDC noise, and compare their recovery using burstegard.
@@
For each waveform, we select $10$ logarithmically spaced distances. For each distance, we inject $100$ signals with random sky position, polarization angle, and cosine of the inclination. A signal is considered to be recovered if there is a trigger within the time and frequency limits of the injection with a FAR lower than $1$ per year, and if there is no trigger that meets these conditions in the data without the injection. Therefore, in the case where an unmodelled injection and a loud CBC event overlap, the unmodelled injection is rejected. We use the fraction of recovered signals at a given distance as an estimator for the detection efficiency. The final figure of merit for a given waveform is the distance at which detection efficiency reaches $50$. These values are reported in Table 1 for both cases.
@@ V Conclusion
We have analyzed a mock data challenge consisting of one month of data simulating a triangular Eintein Telescope and a realistic distribution of CBC signals with PySTAMPAS, a data analysis pipeline to search for unmodelled, long-duration GW signals. The main conclusions of this work are summarized in the following points.
  • Regarding background estimation, the time-slides method tends to greatly overestimate the rate of background triggers, because the fraction of overlapping signals in time-shifted data is no longer negligible compared to second-generation detectors. The null stream provides a convenient way to overcome this issue, but it is only available for specific geometrical configurations such as a triangle, and in practice would be affected by calibration errors and non-Gaussian noise. Future, more realistic MDCs will be necessary to properly assess if the null stream can really be used as a replacement for the time-slide method.
@@
  • PySTAMPAS is able to detect a large fraction of the total BBH in the universe, especially at high masses. Although it is not a dedicated CBC pipeline and therefore remains less sensitive than matched filtering techniques, the comparatively much lower computational cost could make it interesting for low-latency searches and independent confirmation of detections. It should also be more robust to “exotic” CBC signals whose parameters do not fall within the template bank of matched-filtering-based pipelines, such as systems with high eccentricities, matter effects or strong higher-order modes, as well as to uncertainties in the waveforms.
  • Using a dedicated seedless clustering algorithm that tries to fit Newtonian chirps onto the tf-maps, the search is able to recover a larger amount of BNS, up to $z=0.68$, with a detection rate of around $3$ per day, and to estimate the chirp mass and coalescence time with reasonable accuracy ($\sim 1\%$) given the simplicity of the approach. This might be sufficient to trigger full Bayesian inference on the candidates, for more precise and accurate results.
@@
We have therefore shown that an unmodeled search algorithm would be able to detect a significant fraction of CBC signals present in ET, and to provide an accurate estimation of the chirp mass of BNS. The whole search, including the analysis of $102$ years of background, took $\sim 20$ hours using $120$ virtualized off-the-shelf x86_64 CPU cores, which is much lower than what would likely be required by template-based searches Lenon et al. (2021). We stress that this is an analysis of one-month of ET data that can be carried out today with commonplace computing resources. The relatively low computational cost of this search could make it interesting for a quick look of ET data, for early-warning BNS alerts, and for an independent validation of detections in a context where a large number of high-SNR signals are present in the data and may not be easy to model with a sufficient accuracy for deep matched-filter searches.

### v2

@@ IV Effect of the CBC foreground on the detectability of other transient signals
An important issue regarding ET data analysis is the effect of the CBC foreground on the detection capabilities of other GW sources that are expected to be present in the data. In the case of stochastic GW backgrounds, it has been shown that the CBC foreground will indeed prevent the observation of a stochastic background of cosmological origin, which motivates the development of methods to accurately substract CBC signals from the data Harms et al. (2008); Martinovic et al. (2021); Biscoveanu et al. (2020); Sachdev et al. (2020). Here, we study how CBC events recovered by PySTAMPAS affect the ability of the pipeline to detect other kind of unmodeled signals. In order to do that, we inject a selection of waveforms routinely used to characterize unmodeled long-duration GW searches on top of the MDC data, and on top of pure Gaussian noise following the same sensitivity curve as the MDC noise, and compare their recovery using burstegard.
@@
For each waveform, we select $10$ logarithmically spaced distances. For each distance, we inject $100$ signals with random sky position, polarization angle, and cosine of the inclination. A signal is considered to be recovered if there is a trigger within the time and frequency limits of the injection with a FAR lower than $1$ per year, and if there is no trigger that meets these conditions in the data without the injection. Therefore, in the case where an unmodeled injection and a loud CBC event overlap, the unmodeled injection is rejected. We use the fraction of recovered signals at a given distance as an estimator for the detection efficiency. The final figure of merit for a given waveform is the distance at which detection efficiency reaches $50$. These values are reported in Table 1 for both cases.
@@ V Conclusion
We have analyzed a mock data challenge consisting of one month of data simulating a triangular Eintein Telescope and a realistic distribution of CBC signals with PySTAMPAS, a data analysis pipeline to search for unmodeled, long-duration GW signals. The main conclusions of this work are summarized in the following points.
  • Regarding background estimation, the time-slides method tends to greatly overestimate the rate of background triggers, because the fraction of overlapping signals in time-shifted data is no longer negligible compared to second-generation detectors. The null stream provides a convenient way to overcome this issue, but it is only available for specific geometrical configurations such as a triangle, and in practice would be affected by calibration errors and non-Gaussian noise. Future, more realistic MDCs will be necessary to properly assess if the null stream can really be used as a replacement for the time-slide method. In this search, using the null stream to estimate the background instead of the time-slides method allows to recover $86\%$ more signals at a fixed FAR threshold.
@@
  • PySTAMPAS is able to detect a large fraction of the total BBH in the universe, especially at high masses. Although it is not a dedicated CBC pipeline and therefore remains less sensitive than matched filtering techniques, the comparatively much lower computational cost could make it interesting for low-latency searches and independent confirmation of detections. It should also be more robust to “exotic” CBC signals whose parameters do not fall within the template bank of matched-filtering-based pipelines, such as systems with high eccentricities, matter effects or strong higher-order modes, as well as to uncertainties in the waveforms. The better sensitivity of ET at low frequency increases the effective duration of BBH signals, which drives them in the parameter space covered by this search algorithm.
  • Using a dedicated seedless clustering algorithm that tries to fit Newtonian chirps onto the tf-maps, the search is able to recover a larger amount of BNS, up to $z=0.68$, with a detection rate of around $3$ per day, and to estimate the chirp mass and coalescence time with reasonable accuracy ($\sim 1\%$) given the simplicity of the approach. This might be sufficient to trigger full Bayesian inference on the candidates, for more precise and accurate results. This approach was not designed for short duration signals and performs poorly on NSBH and BBH as expected.
@@
We have therefore shown that PySTAMPAS would be able to detect a significant fraction of CBC signals present in ET, and to provide a much quicker estimation of the chirp mass of BNS than modeled pipelines, with an average precision of 1.3% as a byproduct of the search algorithm.
The whole search, including the analysis of $102$ years of background, took $\sim 20$ hours using $120$ virtualized off-the-shelf x86_64 CPU cores, which is much lower than what would likely be required by template-based searches Lenon et al. (2021). We stress that this is an analysis of one-month of ET data that can be carried out today with commonplace computing resources. The relatively low computational cost of this search could make it interesting for a quick look of ET data, for early-warning BNS alerts, and for an independent validation of detections in a context where a large number of high-SNR signals are present in the data and may not be easy to model with a sufficient accuracy for deep matched-filter searches.